OpenStudy (anonymous):

A greenhouse that specializes in growing bell peppers is divided into sections. The number of plants in each section depends on the number of sprinklers in that section. In a section with x sprinklers, there are 3x(x + 3) red bell pepper plants and (x + 4)2 yellow bell pepper plants. Which simplified expression represents the total number of red and yellow bell pepper plants in a section with x sprinklers? 3x2 + 17x + 16 4x2 + 17x + 16 3x4 + 17x2 + 16 4x4 + 17x2 + 16

OpenStudy (ganpat):

It says, The number of plants in each section depends on the number of sprinklers in that section. 3x(x + 3) + (x + 4)2 = x can u simplify further ???

OpenStudy (anonymous):

x=-2?

OpenStudy (ganpat):

3x2 + 6x +2x + 8 = x 3x2 +7x + 8 =0 it is supposed to be 1st option.. but m not sure where does 16 comes from.. My calculations calculates 8..

OpenStudy (campbell_st):

its a case of expanding both expressions the number of red and yellow bell peppers is \[3x(x + 3) + (x + 4)^2\] which expands to \[3x^2 + 9x + x^2 + 8x + 16\] collect like terms for the answer

OpenStudy (ganpat):

ohkey, its squared, not multiplied by 2...